To jump, you look exert a force of 1400 N of force on the ground. (it exerts 1400 Newtons back onto you) You apply this force over a distance of 0.3 m as you extend your legs. How much work have you done? How high will you rise into the air if your mass is 70 kg?
1
Expert's answer
2020-11-25T10:50:50-0500

First, find how stiff your legs are:


"k=\frac{F}{x}."


The work is

"W=\frac12kx^2=\frac{1}{2}Fx=210\text{ J}."

If we spend this work to increase potential energy:


"W=E_P=mgh,"

the height will be


"h=\frac{W}{mg}=0.31\text{ m}."